# Image Inpainting via Tractable Steering of Diffusion Models

### v1

@@ 1 Introduction
Having observed that the lack of tractability hinders us from fully exploiting diffusion models in constrained generation tasks, we study the converse problem: what is the benefit of models that by design support efficient constrained generation? This paper presents positive evidence by showing that Probabilistic Circuits (PCs) (Choi et al., 2020), a class of expressive Tractable Probabilistic Models (TPMs) that support efficient computation of arbitrary marginal probabilities, can efficiently steer the denoising process of diffusion models towards high-quality inpainted images. We will define a class of constraints that includes inpainting constraints for which we can provide the following guarantee. For any constraint $c$ in this class, given a sample $\bm{x}_{t}$ at noise level $t$, we show that a PC trained on noise-free samples (i.e., ${p}(\mathbf{X}_{0})$) can be used to efficiently compute ${p}(\bm{x}_{0}|\bm{x}_{t},c)$, which is a key step in the sampling process of diffusion models. This PC-computed distribution can then be used to effectively guide the denoising process, leading to photorealistic images that adhere to the constraints. Figure 1 illustrates the steering effect of PCs in the proposed algorithm Tiramisu (Tractable Image Inpainting via Steering Diffusion Models). Specifically, we plot the reconstructed image by the diffusion model (the first row of Tiramisu) and the PC (the third row) at five time steps during the denoising process. Compared to the diffusion model baselines, Tiramisu generates more semantically coherent images with the PC-provided guidance. In summary, this paper has three main contributions:
@@ 4.2 Computing Constrained Posterior Distribution
We call $w_{i}$ the soft-evidence constraint of variable $X_{0}^{i}$ as it defines a prior belief of its value. In the extreme case of conditioning on hard evidence, $w_{i}$ becomes an indicator that puts all weight on the conditioned value. Recall from Section 2 that diffusion models parameterize ${p}_{\theta}(\tilde{\bm{x}}_{0}|\bm{x}_{t})$ as a fully-factorized distribution. In order to compute the weighted geometric mean of ${p}_{\mathrm{DM}}(\tilde{\bm{x}}_{0}|\bm{x}_{t},\bm{x}_{0}^{\mathrm{k}})$ and ${p}_{\mathrm{TPM}}(\tilde{\bm{x}}_{0}|\bm{x}_{t},\bm{x}_{0}^{\mathrm{k}})$, we need to also compute the univariate distributions ${p}_{\mathrm{TPM}}(\tilde{x}_{0}^{i}|\bm{x}_{t},\bm{x}_{0}^{\mathrm{k}})$ ($\forall\tilde{X}_{0}^{i}\in\tilde{\mathbf{X}}_{0}$). While this seems to suggest the need to query the PC at least $|\tilde{\mathbf{X}}_{0}|$ times, we propose an algorithm that only needs a forward and a backward pass to compute all target probabilities.
@@
where $\mathsf{pa}(n)$ is the set of parents of node $n$. Next, we gather all input nodes defined on $X_{0}^{i}$ ($\forall i$), denoted $S_{i}$, and compute ${p}_{\mathrm{TPM}}(\tilde{x}_{0}^{i}|\bm{x}_{t},\bm{x}_{0}^{\mathrm{k}}):=\frac{1}{Z}\sum_{n\in S_{i}}\mathtt{bk}_{n}\cdot f_{n}(x_{0}^{i})\cdot w_{i}(x_{0}^{i})$. We justify the correctness of this algorithm in the following theorem, whose proof is provided in Appendix A.
@@
For any smooth and decomposable PC ${p}(\mathbf{X})$ and univariate weight functions $\{w_{i}(X_{i})\}_{i}$, define ${p}^{\prime}(\bm{x})=\frac{1}{Z}\prod_{i}w_{i}(x_{i})\cdot{p}(\bm{x})$. The above-described algorithm computes ${p}^{\prime}(x_{i})$ for all variables $x_{i}$ in time $\mathcal{O}(|{p}|+|\mathbf{X}|)=\mathcal{O}(|{p}|)$.
@@ 6 Experiments
In this section, we take gradual steps to analyze and illustrate our method Tiramisu (Tractable Image Inpainting via Steering Diffusion Models). Specifically, we first qualitatively investigate the steering effect of the TPM on the denoising diffusion process (Sec. 6.1). Next, we perform an empirical evaluation of Tiramisu on three high-resolution image datasets with six large-hole masks, which significantly challenges its ability to generate semantically consistent images (Sec. 6.2). Finally, inspired by the fact that Tiramisu can handle arbitrary constraints that can be written as independent soft evidence (cf. Sec. 4.2), we test it on a new controlled image generation task termed image semantic fusion, where we are given parts (e.g., patches) from different images and aim to fuse them to be both semantically coherent and of high fidelity (Sec. 6.3).
@@ Acknowledgements
To facilitate reproducibility, we have clearly stated the proposed PC inference algorithm in Section 4.2. Algorithmic details including the choice of hyperparameters are provided in Appx. B and C.2. Formal proof of Thm. 1 is provided in Appendix A.

### v2

@@ 1 Introduction
Having observed that the lack of tractability hinders us from fully exploiting diffusion models in constrained generation tasks, we study the converse problem: what is the benefit of models that by design support efficient constrained generation? This paper presents positive evidence by showing that Probabilistic Circuits (PCs) (Choi et al., 2020), a class of expressive Tractable Probabilistic Models that support efficient computation of arbitrary marginal probabilities, can efficiently steer the denoising process of diffusion models towards high-quality inpainted images. We will define a class of constraints that includes inpainting constraints for which we can provide the following guarantee. For any constraint $c$ in this class, given a sample $\bm{x}_{t}$ at noise level $t$, we show that a PC trained on noise-free samples (i.e., ${p}(\mathbf{X}_{0})$) can be used to efficiently compute ${p}(\bm{x}_{0}|\bm{x}_{t},c)$, which is a key step in the sampling process of diffusion models. This PC-computed distribution can then be used to effectively guide the denoising process, leading to photorealistic images that adhere to the constraints. Figure 1 illustrates the steering effect of PCs in the proposed algorithm Tiramisu (Tractable Image Inpainting via Steering Diffusion Models). Specifically, we plot the reconstructed image by the diffusion model (the first row of Tiramisu) and the PC (the third row) at five time steps during the denoising process. Compared to the baselines, Tiramisu generates more semantically coherent images with the PC-provided guidance. In summary, this paper has three main contributions:
@@ 4.2 Computing Constrained Posterior Distribution
We call $w_{i}$ the soft-evidence constraint of variable $X_{0}^{i}$ as it defines a prior belief of its value. In the extreme case of conditioning on hard evidence, $w_{i}$ becomes an indicator that puts all weight on the conditioned value. Recall from Section 2 that diffusion models parameterize ${p}_{\theta}(\tilde{\bm{x}}_{0}|\bm{x}_{t})$ as a fully-factorized distribution. In order to compute the weighted geometric mean of ${p}_{\mathrm{DM}}(\tilde{\bm{x}}_{0}|\bm{x}_{t},\bm{x}_{0}^{\mathrm{k}})$ and ${p}_{\mathrm{TPM}}(\tilde{\bm{x}}_{0}|\bm{x}_{t},\bm{x}_{0}^{\mathrm{k}})$ (cf. Eq. (5)), we need to also compute the univariate distributions ${p}_{\mathrm{TPM}}(\tilde{x}_{0}^{i}|\bm{x}_{t},\bm{x}_{0}^{\mathrm{k}})$ for every $\tilde{X}_{0}^{i}\!\in\!\tilde{\mathbf{X}}_{0}$. While this seems to suggest the need to query the PC at least $|\tilde{\mathbf{X}}_{0}|$ times, we propose an algorithm that only needs a forward and a backward pass to compute all target probabilities.
@@
where $\mathsf{pa}(n)$ is the set of parents of node $n$. Next, for every $i$, we gather all input nodes defined on $X_{0}^{i}$, denoted $S_{i}$, and compute ${p}_{\mathrm{TPM}}(\tilde{x}_{0}^{i}|\bm{x}_{t},\bm{x}_{0}^{\mathrm{k}}):=\frac{1}{Z}\sum_{n\in S_{i}}\mathtt{bk}_{n}\cdot f_{n}(x_{0}^{i})\cdot w_{i}(x_{0}^{i})$. We justify the correctness of this algorithm in the following theorem, whose proof is provided in Appendix A.
@@
For any smooth and decomposable PC ${p}(\mathbf{X})$ and univariate weight functions $\{w_{i}(X_{i})\}_{i}$, define ${p}^{\prime}(\bm{x})=\frac{1}{Z}\prod_{i}w_{i}(x_{i})\cdot{p}(\bm{x})$, where the normalizing constant $Z:=\sum_{\bm{x}}\prod_{i}w_{i}(x_{i})\cdot{p}(\bm{x})$. Assume all variables in $\mathbf{X}$ are categorical variables with $C$ categories, the above-described algorithm computes ${p}^{\prime}(x_{i})$ for every variable $X_{i}$ and its every assignment $x_{i}$ in time $\mathcal{O}(|{p}|+|\mathbf{X}|\cdot C)=\mathcal{O}(|{p}|)$.
@@ 6 Experiments
In this section, we take gradual steps to analyze and illustrate our method Tiramisu (Tractable Image Inpainting via Steering Diffusion Models). Specifically, we first qualitatively investigate the steering effect of the TPM on the denoising diffusion process (Sec. 6.1). Next, we perform an empirical evaluation on three high-resolution image datasets with six large-hole masks, which significantly challenges its ability to generate semantically consistent images (Sec. 6.2). Finally, inspired by the fact that Tiramisu can handle arbitrary constraints that can be written as independent soft evidence (cf. Sec. 4.2), we test it on a new controlled image generation task termed image semantic fusion, where the goal is to fuse parts from different images and generate images with semantic coherence and high fidelity (Sec. 6.3). Code is available at https://github.com/UCLA-StarAI/Tiramisu.
@@ Acknowledgements
To facilitate reproducibility, we have clearly stated the proposed PC inference algorithm in Section 4.2. Algorithmic details including the choice of hyperparameters are provided in Appx. B and C.2. Formal proof of Thm. 1 is provided in Appendix A. We provide the code to train the PCs and to generate inpainted images at https://github.com/UCLA-StarAI/Tiramisu.